The flex-headed ratchet wrench is subjected to a f
Answer #1

The moment of force is given by:

M = Fd

where d is the perpendicular distance:

then d is:

d = 0.75 in+10 in(\sin(60))

then the moment is:

M = 16(0.75+10\sin(60)) = 150 lb. in